3. Goods market equilibrium An economy is described by: C = 160 + 0.6YD I = 150 G = 150 T = 100 NX = 0

a. Find equilibrium Y (real GDP).

b. Find disposable income, YD

c. Consumer spending. Now let G = 200.

d. Find the new value of Y.

e. Find the new value of disposable income.

f. Compute private saving, public saving and national saving.

g. Does national saving = I in this case. Show why or why not.
Answer #1

3. We are given that

C = 160 + 0.6YD I = 150 G = 150 T = 100 NX = 0

a. We know that equilibrium Y (real GDP) has Y = C + I + G + NX

Y = 160 + 0.6*(Y - 100) + 150 + 150 + 0

0.4Y = 400

Y = 1000

Hence equilibrium Y is 1000

b. Disposable income, YD = Y - T = 1000 - 100 = 900

c. Consumer spending C = 160 + 0.60*900 = 700

Now let G = 200.

d. G is increased by 50 and multiplier is 1/(1-MPC) = 1/(1-0.6) = 2.5. Hence Y increases by 50*2.5 = 125. Hence new Y is 1000 + 125 = 1125.

e. New value of disposable income YD = 1125 - 100 = 1025

f. Private saving = YD - C = 1025 - 160 - 1025*0.6 = 250

Public saving = T - G = 100 - 200 = -100

National saving = 250 - 100 = 150

g. National saving = I in this case as both are 150. This is because NX is 0.